OpenStudy (anonymous):

the volume of a sphere is 3,500pie m^3. what is the surface area of the sphere to the nearest square meter

OpenStudy (bahrom7893):

ughh too lazy lol ima go eat breakfast, but this is how u'd do it: V = (4/3) pi r^3 = 3500pi

OpenStudy (bahrom7893):

(4/3)r^3 = 3500, solve for r then find the surface area by pluggin r into the following: A = 4pi r^2

OpenStudy (anonymous):

lol ur too lazy or me? lol i like bein lazy but i jus dnt get volume

OpenStudy (anonymous):

ok thnx

OpenStudy (bahrom7893):

just solve for r and then put that into the surface area..

OpenStudy (anonymous):

how do you plug in for surface area???

OpenStudy (anonymous):

surface area = 4 pi r^2

OpenStudy (anonymous):

LOL, u tryin to slow him down:-)

OpenStudy (anonymous):

ok i got 835.561??????

OpenStudy (anonymous):

this website seems slow this morning what did you get for r?

OpenStudy (anonymous):

umm idk

OpenStudy (anonymous):

r^3 = 3500 / (4pi/3)

OpenStudy (anonymous):

couldnt i jus cube root it enstead of tht confusing division an get 15.18

OpenStudy (anonymous):

= 835.561 r = cube root of this

OpenStudy (anonymous):

so 9.4

OpenStudy (anonymous):

= 9.419 - to get a more accurate result

OpenStudy (anonymous):

noe plug this into the formula for surface area 4 pi r^2

OpenStudy (anonymous):

Surface area is 3*vol/r

OpenStudy (anonymous):

1114.858

OpenStudy (anonymous):

???

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

but thats not one of my answer choices

OpenStudy (anonymous):

another way is to use the formula estudier posted - plug in volume and radius

OpenStudy (anonymous):

what are your answer choices

OpenStudy (anonymous):

a) 2391 m^2 b)43982 m^2 c) 173 m^2 d) 1196 m^2

OpenStudy (anonymous):

ok i'll recheck the calculations

OpenStudy (anonymous):

oh ok - i used 3500 for the volume - it is 3500 pi - sorry

OpenStudy (anonymous):

so??? wats tha answer?

OpenStudy (anonymous):

so we have 4/3 r^3 = 3500 r = 13.795

OpenStudy (anonymous):

oh snap its A

OpenStudy (anonymous):

using the formula are = 3 * vol / r area = 3 * 3500pi / 13.795 = 2391 to nearest meter

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

ok how you get surface area from volume???

OpenStudy (anonymous):

V=4 pi r^3 /3 SA = 4 pi r^2 -> SA = 3*V/r (divide on by other).

OpenStudy (anonymous):

ok thnx imma try an get it myself

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

ok one more ? how you get r from volume?

OpenStudy (anonymous):

rearrange the formula for voluime making r the subject V = (4/3) pi r^3 divide both sides by (4/3) pi r^3 = V / (4/3) pi r = cube root [ V / (4/3) pi]

OpenStudy (anonymous):

In this case you didn't have to worry about the pi because the question put the volume as a multiple of pi, so it cancels.

OpenStudy (anonymous):

ok imma post the ? that im workin on hold on